ERRS continued excavation along the northern slope of the Site. XRF screening data showed concentrations of lead in these excavated soils up to 57,895 ppm lead. Excavation was sloped towards the southern area of the contaminated hillside, moving away from Beaver Creek. A 5-foot concrete wall was unearthed. ERRS continued to unearth battery casings in this area and lead concentrations were determined up to 65,023 ppm. START continued to screen the areas beneath excavation to determine "clean" areas with lead concentrations below 1,000 ppm. START collected samples from approximately 20% of these locations and shipped the samples for laboratory analysis and confirmation. START continued to draft maps that contain Site photographs of the excavation areas depicting XRF data at the sample locations.
ERRS received analytical data from Stockpile #2. Lead TCLP was 48.9 mg/L; the soil was designated "hazardous". The estimated volume of soil in both Stockpiles #1 and #2 was 2,800 tons.
ERRS received bids for T&D of the stockpiled lead-contaminated soil and awarded the subcontract to the lowest bidder. ERRS began T&D operations on February 3, 2010. A total of 30 trucks, estimated at 22 tons each, were loaded with hazardous soil from stockpile #1 through February 4, 2010.
T&D operations continued through February 12, 2010, during days that did not pose unfavorable weather conditions. During the week of February 8, 2010, a total of 80 trucks, carrying approximately 22 tons each, were loaded with hazardous soil. ERRS completed loading soil from Stockpile #1 and began loading soil from Stockpile #2. To date, 110 trucks, approximately 2,420 tons, were loaded with hazardous soil. The soil was transported to Detroit, Michigan, for disposal. Approximately 400 tons of soil remain stockpiled on Site, awaiting T&D.
ERRS continued to stockpile fill material on Site, in preparation for construction of a cap over the remaining contaminated soils. ERRS are utilizing a large pile of fill material that was already present on the Site.
START continued to conduct air monitoring with DataRAM 4000's along the southern, western, and eastern areas of the Site. The maximum concentration of particulates that migrated from the Site during this operational period was 27.3 ug/m3.
START continued to conduct personal air monitoring for particulates in the hot zones. The instrument was rotated among the operators in the excavator and articulate dump in the excavation area and the laborers and operator in the stockpile area. The maximum concentration of particulates in the hot zone during this operational period was 14 ug/m3. No respiratory protection was necessary during this operational period.
Representatives from VDEQ were on site to observe operations and the condition of the Site. The OSC updated VDEQ on Site operations.
Interested parties may view added recent pictures in the Images section of the website: www.epaosc.org/twincities.
|